A Ferris wheel that can carry almost 100 people has a diameter of 20.0 m. If possible, we

created an electromagnet with 100 coils that can produce a field of an average strength of 0.

90 T parallel to the axis of the wheel. If the field steadily decreases to zero to produce an

induced emf is 50 V, what is the time needed for this induction to happen?


1
Expert's answer
2021-07-02T09:33:32-0400

Gives

Radius (r)=10m

Coil turn (n)=100

Magnetic field (B)=90T

emf(e)=50V

"\theta=0°"

"\phi=B.A"

"\phi=BAcos\theta"

"\phi=BA"

"e=n\frac{d\phi}{dt}"

"A=\pi r^2=3.14\times10^2=314m^2"

"dt=\frac{nd\phi}{e}"

Put value

"dt=\frac{100\times(90-0)\times314 }{50}=56520sec"

"dt=15.7h"